# Adaptive Accountability in Networked MAS:

### v1

@@ Abstract
Large-scale networked multi-agent systems increasingly underpin critical infrastructure, yet their collective behavior can drift toward undesirable emergent norms that elude conventional governance mechanisms. We introduce an adaptive accountability framework that (i) continuously traces responsibility flows through a lifecycle-aware audit ledger, (ii) detects harmful emergent norms online via decentralized sequential hypothesis tests, and (iii) deploys local policy and reward-shaping interventions that realign agents with system-level objectives in near real time. We prove a bounded-compromise theorem showing that whenever the expected intervention cost exceeds an adversary’s payoff, the long-run proportion of compromised interactions is bounded by a constant strictly less than one. Extensive high-performance simulations with up to 100 heterogeneous agents, partial observability, and stochastic communication graphs show that our framework prevents collusion and resource hoarding in at least $90\%$ of configurations, boosts average collective reward by $12\text{–}18\%$, and lowers the Gini inequality index by up to $33\%$ relative to a PPO baseline. These results demonstrate that a theoretically principled accountability layer can induce ethically aligned, self-regulating behavior in complex MAS without sacrificing performance or scalability.
Keywords Multi-Agent $\cdot$ Accountability $\cdot$ Emergent Norms $\cdot$ Norm Detection $\cdot$ Adaptive Governance $\cdot$ Ethical AI
@@ 1 Introduction
Multi-agent systems (MAS) increasingly underpin critical applications in transportation, smart-grid energy management, finance, and healthcare [shoham2007if, wooldridge2009introduction]. By distributing decision making across many partly autonomous agents, these systems offer scalability, resilience to single-point failures, and rapid adaptation to non-stationary environments. Those benefits, however, come with interaction complexity: once deployed, the collective behavior can drift toward undesirable emergent norms that were never explicitly designed or anticipated [stone2000multiagent, Rahwan2019MachineBehavior]. Identifying and correcting such norms is essential for ensuring fairness, security, and compliance with societal values.
In large-scale networked MAS, responsibility is diffused across agents and time steps. Classical AI-accountability frameworks—often built for single models or clearly identifiable human decision makers—rely on static compliance checks or post-hoc audits [mittelstadt2019principles, jobin2019global]. These approaches break down when no single node sees the full global state and harms can emerge from subtle feedback loops. Recent policy instruments, such as the EU Artificial Intelligence Act [eu_ai_act_2024] and the U.S. NIST AI Risk-Management Framework [nist_ai_rmf_2023], call for “collective accountability” in AI infrastructures but leave open the technical challenge of tracing and mitigating harmful norms online and in a decentralized fashion.
Three established research threads touch on this challenge yet remain individually insufficient. Normative MAS studies formal norms and sanctions [Andrighetto2013NormDynamics, Hollander2011NormativeSystems] but typically assumes a central monitor with perfect state access. Multi-agent reinforcement learning (MARL) has produced sophisticated training algorithms [busoniu2008comprehensive, Foerster2018COMA], but only recently has begun to explore equity, collusion, or convergence to harmful equilibria. Runtime assurance for cyber-physical systems proposes supervisory safety guards [perez2024artificial], yet seldom scales beyond a handful of agents or supports changing objectives. Consequently, the field still lacks a comprehensive method to detect, trace, and correct undesirable emergent norms under partial observability and heterogeneous incentives.
@@
To evaluate the framework at realistic scale, we run extensive simulations on a high-performance cluster equipped with NVIDIA A40 GPUs. Up to 100 heterogeneous agents interact over stochastic communication graphs under partial observability. Across 360 Monte-Carlo configurations that vary random seeds, payoff functions, and intervention protocols, the framework averts collusion and resource hoarding in at least $90\%$ of runs, lifts average collective reward by $12\text{–}18\%$, and reduces the Gini inequality index by up to $33\%$ relative to a PPO baseline. All code and simulation scripts will be released to support replication.
@@ 2 Background and Literature Review
Multi-agent systems (MAS) consist of multiple autonomous entities that sense, decide, and act on local information while exchanging limited signals with neighbors [wooldridge2009introduction, shoham2007if]. Delegating control in this way yields scalability, fault tolerance, and graceful adaptation, enabling real-world deployments in smart-grid scheduling [khan2019multiagent], adaptive traffic control [yan2020reinforcement], supply-chain optimization [giannakis2020multi], cooperative robotics [brambilla2013swarm], distributed optimization on sensor networks [yang2019survey], algorithmic trading [tesfatsion2006agent], and cross-hospital resource sharing [herrouz2019multi]. What unites these domains is the need to reconcile local incentives with global constraints under partial observability and stringent latency budgets.
Nonetheless, the same decentralization that confers robustness breeds interaction complexity. Agents are autonomous, social, and both reactive and proactive [wooldridge1995intelligent]; even simple local rules can induce system-level phenomena that are hard to predict, verify, or control [stone2000multiagent]. Coordination grows more challenging as populations scale, links fluctuate, or objectives diverge [corkill2015collaborating]. Emergent behaviors—benign or harmful—complicate formal safety guarantees in partially observable settings [busoniu2008comprehensive].
@@ 2.1 Emergent Norms: From Descriptive Theory to Algorithmic Design
Classical emergence studies viewed norms as regularities not explicit in any single agent’s code [bedau2002downward]. Early mechanisms included imitation and social learning [sen2007emergence], reinforcement-driven sanctioning [mohammad2021multi], and top-down institutions with explicit penalties [conte2012normative, pitt2012axiomatization]. Recent work shifts from descriptive to constructive: Tzeng2024NormEnforcement show that “soft-touch” speech acts (tell, hint) enforce norms faster than pure sanction; Oldenburg2024SharedNorms introduce Bayesian rule-induction so newcomers infer latent norms on-line; Serramia2024Consensus formalise optimal consensus norms that balance heterogeneous stakeholder utilities; Ren2024CRSEC harness large language models to generate conflict-reducing social norms in an artificial town; and Woodgate2025RawlsianNorms operationalization Rawls’ maximin principle, steering emergent norms toward lower inequality. Together these papers recast norm formation as an algorithmic and ethical design problem.
@@ 2.2 Distributed Accountability and Auditing
Standard artifacts—model cards [mitchell2019model] and fact sheets [arnold2019factsheets]—target monolithic models. Applied to MAS they fail to address diffuse responsibility [rastogi2022distributed], continual policy drift [jobin2019global], and aggregation of local harms. Recent research begins to fill the gap: Chan2024Visibility catalogue visibility tools for peer-to-peer deployments, trading off informativeness and privacy; Mu2024Responsibility extend alternating-time temporal logic with quantitative responsibility metrics; Gyevnar2024CEMA generate natural-language causal explanations via counterfactual roll-outs, boosting trust in autonomous vehicles; and Chang2025Chronicles treat provenance as a chain of custody for LLM-generated artifacts, reconstructable without privileged logs.
@@ 2.3 Intervention and Governance Mechanisms
Detection alone is insufficient; an accountability layer must intervene. Beyond classic reward shaping, recent levers include action-space restriction [Oesterle2024RAISE], monitored MDPs with human oracles [Parisi2024MonitoredMDP], and sample-efficient opponent shaping [Fung2024OpponentShaping]. Model-based credit assignment with counterfactual imagination [Chai2024MACD] shows promise for scalable cooperation—although its computational cost remains high in large MAS.
@@ 2.4 Causal Tracing and Lifecycle Provenance
Effective intervention presupposes causal insight. Triantafyllou2024ASE quantify agent-specific effects—how one agent’s actions propagate to others—while time-uniform concentration bounds [howard2020time] and stochastic approximation theory [kushner2003stochastic] provide rigorous guarantees for sequential detectors and adaptive thresholds such as those used in our framework.
@@ 2.6 Positioning of the Present Work
The Adaptive Accountability Framework introduced next addresses all three needs. It combines a cryptographically grounded Merkle ledger, time-uniform norm detectors, and cost-bounded interventions validated at 100-agent scale. The next sections detail the architecture (Section 4), furnish theoretical guarantees (Section 5), and supply empirical evidence that the framework curbs harmful emergence even under adversarial disturbance (Section 7).
@@ 3 Problem Definition and Research Questions
3.1 Formalising Accountability in Networked MAS
@@
The global state space $\mathcal{S}$ may include physical variables (e.g. queue lengths, robot poses) and latent coordination artefacts (e.g. shared plans). At each round $t\in\mathbb{N}$ the environment occupies $s_{t}\!\in\!\mathcal{S}$. Agent $A_{i}$ acquires a private observation $o_{i,t}\!\in\!\Omega_{i}$ parameterised by a perceptual kernel $P(o_{i,t}\!\mid s_{t},i)$; selects an action $a_{i,t}\!\in\!\mathcal{A}_{i}$ via a history-dependent policy $\pi_{i,t}\!\in\!\Delta(\mathcal{A}_{i})^{\Omega_{i}^{t+1}}$; and receives an individual reward
@@
where $\mathbf{a}_{t}=(a_{1,t},\dots,a_{N,t})$ and $\lambda\!\in\![0,1]$ trades off egoistic and collective incentives. The transition kernel $T$ yields $s_{t+1}\!\sim\!T(s_{t},\mathbf{a}_{t})$; consequently, learning updates $\pi_{i,t+1}\!\leftarrow\!\mathrm{Learn}(\pi_{i,t},\mathcal{D}_{i,t})$ on local traces $\mathcal{D}_{i,t}$, imparting non-stationarity to the joint dynamics $T\!\circ\!\Pi_{t}$ [hernandez2019survey].
@@
where each vertex $e\in\mathcal{E}_{t}$ is an atomic event $\langle\textsf{id},\,t,\,\mathrm{type},\,\mathrm{payload}\rangle$ (actions, messages, external shocks). An edge $(e_{k}\!\rightarrow\!e_{\ell})\!\in\!\mathcal{C}_{T}$ encodes empirical causal influence. Because full interventional data are unavailable online, we approximate causality with rolling multivariate Granger tests on an $h$-step horizon and retain only edges whose F-statistic exceeds a tunable significance level $\alpha_{\!\text{GC}}$; this yields $|\mathcal{C}_{T}|=\mathcal{O}(hN\deg^{+}(G_{t}))$ and keeps storage linear in $T$ [sherman1950adjustment].
@@
where $\phi(e,\mathcal{S})$ equals $1$ if $e$ would still occur when only agents in $\mathcal{S}$ are allowed to influence upstream events and 0 otherwise. We compute $\rho_{i}(e)$ via a Monte-Carlo kernel-SHAP proxy with at most $K$ coalition samples per event, keeping amortised overhead sub-quadratic in $N$.
@@
Threat and intervention model:
Up to $\kappa$ agents are Byzantine: they may deviate arbitrarily from $\pi_{i,t}$, falsify messages, or fork the ledger. The remaining agents are rational learners whose objective coefficients $\lambda$ need not align with social welfare, creating the possibility of “lawful but awful” outcomes. A supervisor (possibly distributed) deploys interventions $\sigma_{t}=\langle\mathcal{I}_{t},\textsf{type},\theta_{t}\rangle$ with
@@ 3.2 Research Questions
Design an event ledger $\mathcal{L}_{T}$ and allocation rule $\rho_{i}(e)$ that remain accurate under partial observability, message loss, and non-stationary policies, yet scale to $N\!\sim\!10^{2}$$10^{3}$ agents [rastogi2022distributed].
@@
Develop sequential tests that flag emergent negative norms with low false-alarm rate using only streaming, decentralised data, achieving $\mathcal{O}(\log T)$ detection delay for i.i.d. deviations and acceptable power under temporal correlation [sen2007emergence].
@@
Identify local interventions $\sigma_{t}$ that minimise compromise $C_{T}$ and welfare loss $\Delta J_{\text{soc}}$ subject to budget $c_{\max}$; analyse stability for reward shaping, model-patch injection, and network throttling [mohammad2021multi].
@@
Assess whether the above mechanisms retain effectiveness as $N$, graph topology, objective heterogeneity, or adversarial mix vary; empirical tests range from $|V|=10$ to 100 and span Watts–Strogatz and scale-free networks.

### v2

@@ Abstract
Large-scale networked multi-agent systems increasingly underpin critical infrastructure, yet their collective behavior can drift toward undesirable emergent norms such as collusion, resource hoarding, and implicit unfairness. We present the Adaptive Accountability Framework (AAF), an end-to-end runtime layer that (i) records cryptographically verifiable interaction provenance, (ii) detects distributional change points in streaming traces, (iii) attributes responsibility via a causal influence graph, and (iv) applies cost-bounded interventions—reward shaping and targeted policy patching—to steer the system back toward compliant behavior. We establish a bounded-compromise guarantee: if the expected cost of intervention exceeds an adversary’s expected payoff, the long-run fraction of compromised interactions converges to a value strictly below one. We evaluate AAF in a large-scale factorial simulation suite (87,48087{,}480 runs across two tasks; up to 100 agents plus a 500-agent scaling sweep; full and partial observability; Byzantine rates up to $10\%$; 10 seeds per regime). Across 324 regimes, AAF lowers the executed compromise ratio relative to a Proximal Policy Optimization baseline in 96% of regimes (median relative reduction 11.9%) while preserving social welfare (median change 0.4%). Under adversarial injections, AAF detects norm violations with a median delay of 71 steps (interquartile range 39–177) and achieves a mean top-ranked attribution accuracy of 0.97 at 10% Byzantine rate.
Keywords multi-agent systems $\cdot$ AI accountability $\cdot$ ethical AI $\cdot$ security $\cdot$ governance $\cdot$ reinforcement learning
@@ 1 Introduction
Multi-agent systems (MAS) increasingly underpin critical applications in transportation, smart-grid energy management, finance, and healthcare [59, 68]. By distributing decision making across many partly autonomous agents, these systems offer scalability, resilience to single-point failures, and rapid adaptation to non-stationary environments. Those benefits, however, come with interaction complexity: once deployed, the collective behavior can drift toward undesirable emergent norms that were never explicitly designed or anticipated [61, 52]. Identifying and correcting such norms is essential for ensuring fairness, security, and compliance with societal values.
In large-scale networked MAS, responsibility is diffused across agents and time steps. Classical AI-accountability frameworks—often built for single models or clearly identifiable human decision makers—rely on static compliance checks or post-hoc audits [38, 28]. These approaches break down when no single node sees the full global state and harms can emerge from subtle feedback loops. Recent policy instruments, such as the EU Artificial Intelligence Act [17] and the U.S. NIST AI Risk-Management Framework [42], call for “collective accountability” in AI infrastructures but leave open the technical challenge of tracing and mitigating harmful norms online and in a decentralized fashion.
Three established research threads touch on this challenge yet remain individually insufficient. Normative MAS studies formal norms and sanctions [4, 25] but typically assumes a central monitor with perfect state access. Multi-agent reinforcement learning (MARL) has produced sophisticated training algorithms [9, 19], but only recently has begun to explore equity, collusion, or convergence to harmful equilibria. Runtime assurance for cyber-physical systems proposes supervisory safety guards [49], yet seldom scales beyond a handful of agents or supports changing objectives. Consequently, the field still lacks a comprehensive method to detect, trace, and correct undesirable emergent norms under partial observability and heterogeneous incentives.
@@
To evaluate the framework at realistic scale, we run a Q1-scale factorial simulation suite comprising 87,48087{,}480 episodes (2 tasks; $N\leq 100$ (plus a scaling sweep to $N=500$ in resource_sharing); full and partial observability; Byzantine rates up to $10\%$; 10 seeds per regime). Across 324 regimes, AAF reduces the executed compromise ratio relative to PPO-only in 96% of regimes (median reduction 11.9%; $p<0.05$ in 86.4%) while preserving social welfare (median change 0.4%). Inequality effects are mixed, highlighting a tunable compromise–fairness trade-off (see §7.3). All code and simulation scripts will be released to support replication.
@@ 2 Background and Literature Review
Multi-agent systems consist of multiple autonomous entities that sense, decide, and act on local information while exchanging limited signals with neighbors [68, 59]. Delegating control in this way yields scalability, fault tolerance, and graceful adaptation, enabling real-world deployments in smart-grid scheduling [2], adaptive traffic control [51], supply-chain optimization [41], cooperative robotics [8], distributed optimization on sensor networks [14], algorithmic trading [62], and cross-hospital resource sharing [48]. What unites these domains is the need to reconcile local incentives with global constraints under partial observability and stringent latency budgets.
Nonetheless, the same decentralization that confers robustness breeds interaction complexity. Agents are autonomous, social, and both reactive and proactive [67]; even simple local rules can induce system-level phenomena that are hard to predict, verify, or control [61]. Coordination grows more challenging as populations scale, links fluctuate, or objectives diverge [13]. Emergent behaviors—benign or harmful—complicate formal safety guarantees in partially observable settings [9].
@@ 2.1 Emergent Norms: From Descriptive Theory to Algorithmic Design
Classical emergence studies viewed norms as regularities not explicit in any single agent’s code [7]. Early mechanisms included imitation and social learning [56], reinforcement-driven sanctioning [39], and top-down institutions with explicit penalties [5, 50]. Recent work shifts from descriptive to constructive: Tzeng et al. [64] show that “soft-touch” speech acts (tell, hint) enforce norms faster than pure sanction; Oldenburg and Zhi-Xuan [45] introduce Bayesian rule-induction so newcomers infer latent norms on-line; Serramia et al. [57] formalize optimal consensus norms that balance heterogeneous stakeholder utilities; Ren et al. [54] harness large language models to generate conflict-reducing social norms in an artificial town; and Woodgate et al. [66] operationalization Rawls’ maximin principle, steering emergent norms toward lower inequality. Together these papers recast norm formation as an algorithmic and ethical design problem.
@@ 2.2 Distributed Accountability and Auditing
Standard artifacts—model cards [37] and fact sheets [6]—target monolithic models. Applied to MAS they fail to address diffuse responsibility [18], continual policy drift [28], and aggregation of local harms. The recent explosion of Large Language Model (LLM)-based multi-agent systems [33] has further exacerbated these challenges, as generative agents exhibit complex emergent behaviors that are difficult to trace.
Recent research begins to fill this gap: Chan et al. [11] catalogue visibility tools for peer-to-peer deployments, trading off informativeness and privacy; Mu and Oren [40] extend alternating-time temporal logic with quantitative responsibility metrics; Gyevnar et al. [22] generate natural-language causal explanations via counterfactual roll-outs, boosting trust in autonomous vehicles; and Chang and Echizen [12] treat provenance as a chain of custody for LLM-generated artifacts. Ebrahimi and Asudeh [15] systematically survey LLM-based MAS through the “pillars of responsibility” (reliability, transparency, accountability, and fairness), while Raza et al. [53] adapt the Trust, Risk, and Security Management (TRiSM) framework specifically for agentic AI. Furthermore, Solomon and others [60] propose LumiMAS, a comprehensive framework for real-time monitoring and enhanced observability in multi-agent systems, highlighting the critical need for runtime oversight.
@@ 2.3 Intervention and Governance Mechanisms
Detection alone is insufficient; an accountability layer must intervene. Beyond classic reward shaping, recent levers include action-space restriction [44], monitored MDPs with human oracles [47], and sample-efficient opponent shaping [20]. In the realm of safe MARL, constrained policy optimization and Lagrangian relaxation are widely used to enforce safety constraints [1]. Similarly, Li et al. [32] explore Byzantine robust cooperative MARL, formulating the problem as a Bayesian game to maintain coordination even when a fraction of agents act adversarially. Zheng and others [69] further rethink MAS reliability from the perspective of Byzantine fault tolerance, emphasizing the need for robust consensus mechanisms in agentic systems. Model-based credit assignment with counterfactual imagination [10] shows promise for scalable cooperation—although its computational cost remains high in large MAS.
@@ 2.4 Causal Tracing and Lifecycle Provenance
Effective intervention presupposes causal insight. Counterfactual credit assignment has long been used in MARL (e.g., COMA) [19], and more recent work extends causal effect decomposition to multi-agent sequential decision making [63]. In offline settings, Wang and others [65] propose MACCA, a framework that assigns credit by analyzing causal relationships in individual rewards. Time-uniform concentration bounds [26] and stochastic approximation theory [31] provide rigorous guarantees for sequential detectors and adaptive thresholds such as those used in our framework.
@@ 2.6 Positioning of the Present Work
The Adaptive Accountability Framework introduced next addresses all three needs. It combines a cryptographically grounded Merkle ledger, time-uniform norm detectors, and cost-bounded interventions validated at 100-agent scale (and up to 500 agents in a scaling sweep). The next sections detail the architecture (Section 4), furnish theoretical guarantees (Section 5), and supply empirical evidence that the framework curbs harmful emergence even under adversarial disturbance (Section 7).
@@ 3 Problem Definition and Research Questions
3.1 Formalizing Accountability in Networked Multi-Agent Systems
@@
The global state space $\mathcal{S}$ may include physical variables (e.g. queue lengths, robot poses) and latent coordination artifacts (e.g. shared plans). At each round $t\in\mathbb{N}$ the environment occupies $s_{t}\!\in\!\mathcal{S}$. Agent $A_{i}$ acquires a private observation $o_{i,t}\!\in\!\Omega_{i}$ parameterized by a perceptual kernel $P(o_{i,t}\!\mid s_{t},i)$; selects an action $a_{i,t}\!\in\!\mathcal{A}_{i}$ via a history-dependent policy $\pi_{i,t}\!\in\!\Delta(\mathcal{A}_{i})^{\Omega_{i}^{t+1}}$; and receives an individual reward
@@
where $\mathbf{a}_{t}=(a_{1,t},\dots,a_{N,t})$ and $\lambda\!\in\![0,1]$ trade off egoistic and collective incentives. The transition kernel $T$ yields $s_{t+1}\!\sim\!T(s_{t},\mathbf{a}_{t})$; consequently, learning updates $\pi_{i,t+1}\!\leftarrow\!\mathrm{Learn}(\pi_{i,t},\mathcal{D}_{i,t})$ on local traces $\mathcal{D}_{i,t}$, imparting non-stationarity to the joint dynamics $T\!\circ\!\Pi_{t}$ [24].
@@
where each vertex $e\in\mathcal{E}_{t}$ is an atomic event $\langle\textsf{id},\,t,\,\mathrm{type},\,\mathrm{payload}\rangle$ (actions, messages, external shocks). An edge $(e_{k}\!\rightarrow\!e_{\ell})\!\in\!\mathcal{C}_{T}$ encodes empirical causal influence. Because full interventional data are unavailable online, we approximate causality with rolling multivariate Granger tests on an $h$-step horizon and retain only edges whose F-statistic exceeds a tunable significance level $\alpha_{\!\text{GC}}$; this yields $|\mathcal{C}_{T}|=\mathcal{O}(hN\deg^{+}(G_{t}))$ and keeps storage linear in $T$ [58].
@@
where $\phi(e,\mathcal{S})$ equals $1$ if $e$ would still occur when only agents in $\mathcal{S}$ are allowed to influence upstream events and 0 otherwise. We compute $\rho_{i}(e)$ via a Monte-Carlo kernel-SHAP proxy with at most $K$ coalition samples per event, keeping amortized overhead sub-quadratic in $N$.
@@
Explicit Threat Model:
We consider a networked MAS where a subset of agents may act adversarially or selfishly, threatening system-level norms. Specifically, we assume:
  1. 1.
    Byzantine Agents: Up to $\kappa$ agents are Byzantine. They possess full knowledge of the system’s state and the accountability framework’s detection thresholds. They may deviate arbitrarily from their nominal policies $\pi_{i,t}$, collude to trigger false alarms, or attempt to stay just below the CUSUM detection threshold (stealthy attacks). However, they cannot break the cryptographic primitives securing the ledger.
  2. 2.
    Rational Learners: The remaining $N-\kappa$ agents are rational reward-maximizers. Their individual objective coefficients $\lambda$ need not align with social welfare, creating the possibility of “lawful but awful” emergent behaviors (e.g., tragedy of the commons) even without malicious intent.
  3. 3.
    Network Adversary: The communication substrate may drop up to a fraction $p_{\text{drop}}$ of messages or reorder them, but it cannot forge cryptographic signatures or permanently partition the network. The graph $G_{t}$ must remain $(\kappa+1)$-vertex-connected on average.
Intervention model:
A supervisor (possibly distributed) deploys interventions $\sigma_{t}=\langle\mathcal{I}_{t},\textsf{type},\theta_{t}\rangle$ with
@@ 3.2 Research Questions
Design an event ledger $\mathcal{L}_{T}$ and allocation rule $\rho_{i}(e)$ that remain accurate under partial observability, message loss, and non-stationary policies, yet scale to $N\!\sim\!10^{2}$$10^{3}$ agents [18].
@@
Develop sequential tests that flag emergent negative norms with low false-alarm rate using only streaming, decentralized data, achieving $\mathcal{O}(\log T)$ detection delay for i.i.d. deviations and acceptable power under temporal correlation [56].
@@
Identify local interventions $\sigma_{t}$ that minimize compromise $C_{T}$ and welfare loss $\Delta J_{\text{soc}}$ subject to budget $c_{\max}$; analyze stability for reward shaping, model-patch injection, and network throttling [39].
@@
Assess whether the above mechanisms retain effectiveness as $N$, objective heterogeneity, observability, or adversarial mix vary; our main factorial grid spans $|V|\in\{10,50,100\}$ and we additionally run a scaling sweep up to $|V|=500$.